Three consecutive integers are such that the first integer plus twice the second integer plus five less than the third integer is 27. What are the integers?
Answer by addingup(3677)

x+2(x+1)+x+2-5 = 27
x+2x+2+x-3 = 27
group like terms (it doesn't change the result, but it's better/easier)
2x+x+x+2-3 = 27
4x-1 = 27
4x = 28
x = 7
-------------------------------
Check:
7+2(8)+(9-5) = 27
7+16+4 = 27 Correct.
:
John